OpenStudy (eric_d):

s=u+1/2at^2......2 s=1/2(u+v)t.......3 v^2=u^2+2as....4 @Somy

OpenStudy (somy):

so i have to get these formulas?

OpenStudy (eric_d):

yes

OpenStudy (eric_d):

The first 1 is like this: From a=v-u/t at=v-u v=u+at........1

OpenStudy (somy):

ok gimme let me try

OpenStudy (somy):

that's too hard hhh

OpenStudy (eric_d):

you did not learn this before?

OpenStudy (somy):

i do know these formulas but we did not learn derivation

OpenStudy (eric_d):

Okay,nvm..

OpenStudy (somy):

sorry T_T

OpenStudy (eric_d):

It's okay I'll try to figure it out....

OpenStudy (somy):

oh lol found it

OpenStudy (somy):

loool ok

OpenStudy (somy):

(v+u)/2 --- average velocity formula S= v*t thus \[s= \frac{ v+u }{ 2 }\times t\]

OpenStudy (somy):

that's for s=1/2(u+v)t.......3

OpenStudy (somy):

now for this s= ut+ 1/2at^2

OpenStudy (somy):

\[v= u + at\] \[s= \frac{ u+v }{ 2 }\times t\] we use these 2 formulas now we substitute v in second equation by first equation \[s= (\frac{ u+u+at }{2}) \times t\] now simplify \[s= \frac{ 2ut }{ 2 } +\frac{ at^2 }{ 2 }\] thus \[s= ut +\frac{ 1 }{ 2 } at^2\]

OpenStudy (somy):

now next one \[v= u+at \] \[s= \frac{ u+v }{ 2 } \times t\] again we use these formulas lets make t as subject in first equation \[t= \frac{ v-u }{ a }\] now lets substitute t in \[s= \frac{ u+v }{ 2 } \times t\] by the t we just made as subject \[s= \frac{ u+v }{ 2 } \times \frac{ v-u }{ a }\] rearranging this \[2as= (u+v)(v-u) \] \[2as= v^2-u^2 \] make v^2 as subject \[v^2= u^2 +2as\]

OpenStudy (somy):

done :) any questions?

OpenStudy (eric_d):

Nope

OpenStudy (eric_d):

Thanks @Somy

OpenStudy (somy):

you are welcome :)))